HETAV PATEL

Grade 11,

In a Young’s double slit experiment, two narrow vertical slits placed 0.800 mm apart are illuminated by the same source of yellow light of wavelength 589 nm. How far are the adjacent bright bands in the interference pattern observed on a screen 2.00 m away?

In a Young’s double slit experiment, two narrow vertical slits placed 0.800 mm apart are illuminated by the same source of yellow light of wavelength 589 nm. How far are the adjacent bright bands in the interference pattern observed on a screen 2.00 m away?

Grade:upto college level

1 Answers

Deepak Patra
9 years ago
Sol. Given that, d = 0.8 mm = 0.8 * 10^–3 m, λ = 589 nm = 589 * 10^–9 m and D = 2 m. So, β = Dλ/d = 589 * 10^-9/0.8 * 10^-3 = 1.47 * 10^-3 m = 147 mm.